### v2

@@ 4 Convergence to steady-state
Let $R:=d(l^{\circ},I_{ss})$. The potential $\Phi$ is Lipschitz continuous; denote the Lipschitz constant by $C_{\Phi}$. Then $\mathsf{OPT}-\Phi(0)\leq C_{\Phi}R$. Let $T_{1}=C_{1}(\mathsf{OPT}-\Phi(0))$, with $C_{1}$ as guaranteed by Lemma 4.12, so that $\ell$ has the weak steady-state property from time $T_{1}$ onwards. Then $T_{1}\leq C_{1}^{\prime}R$, where $C_{1}^{\prime}=C_{1}C_{\Phi}$.
By Lemma 4.14, $\ell$ reaches steady state by time $T_{1}+T_{2}$, where $T_{2}=\Delta\cdot\max_{v\in V}{\gamma_{v}(T_{1})}$. To bound $T_{2}$, we observe that $\gamma_{v}$, viewed as a function of the labeling, is Lipschitz continuous with Lipschitz constant $2|V|$, since the slacks are Lipschitz continuous with Lipschitz constant $2$. Thus $\gamma_{v}(0)\leq 2R|V|$. Since $\|\ell(T_{1})-\ell(0)\|\leq\kappa T_{1}$, this time by Lipschitz continuity of the trajectory, we deduce the very crude bound $\gamma_{v}(T_{1})\leq 2|V|(R+\kappa T_{1})$ for all $v$.
@@ 5 Uniqueness and continuity
Since there exists $\delta>0$ such that for any equilibrium trajectory $\ell$ starting within $B_{\delta}(l^{\circ})$, $E^{*}_{l^{\circ}}\subseteq E^{*}_{\ell(\theta)}$ and $E^{\prime}_{\ell(\theta)}\subseteq E^{\prime}_{l^{\circ}}$ for all $\theta\in[0,\eta]$, we can focus on the local network $\hat{G}$, where only currently active arcs are present and currently resetting arcs become free arcs. So set $\hat{E}\coloneqq E^{\prime}_{l^{\circ}}$ and $\hat{E}^{\infty}\coloneqq E^{*}_{l^{\circ}}$, and define the local network to be $\hat{G}=(V,\hat{E})$ with free arcs $\hat{E}^{\infty}$. Observe that the set of feasible labelings of the new and old generalized network coincide on $B_{\delta}(l^{\circ})$.
@@
We are now ready to prove uniqueness of equilibrium trajectories. The following theorem applies to generalized networks; Theorem 3.1 is the special case $E^{\infty}=\emptyset$.
Theorem 5.2.
Given a generalized network $G=(V,E)$ with free arcs $E^{\infty}\subseteq E$, for any feasible labeling $l^{\circ}$, there is a unique equilibrium trajectory $\ell$ with $\ell(0)=l^{\circ}$.
Proof.
@@
Let $\theta_{1},\theta_{2},\ldots$ be any sequence of strictly positive times converging to zero. For each $i$, we can consider $\theta\mapsto\ell^{(1)}(\theta+\theta_{i})$ as an equilibrium trajectory starting from $\ell^{(1)}(\theta_{i})$, and $\theta\mapsto\ell^{(2)}(\theta+\theta_{i})$ as an equilibrium trajectory starting from $\ell^{(2)}(\theta_{i})$. By Lemma 5.1, there exists an $\eta>0$ so that $\theta\mapsto\ell^{(1)}(\theta+\theta_{i})$ converges to $\theta\mapsto l^{\circ}+(\theta+\theta_{i})X(l^{\circ})$ on the interval $[0,\eta]$. Taking a limit as $i\to\infty$, we find that $\ell^{(1)}$ must be identical to $\theta\mapsto l^{\circ}+\theta X(l^{\circ})$ on the interval $[0,\eta]$. But the identical argument applies to $\ell^{(2)}$, contradicting our assumption that the trajectories $\ell^{(1)}$ and $\ell^{(2)}$ diverge at time 0. ∎
@@
Also in this section we state and prove a slightly more general version of Theorem 3.2.
Theorem 5.3.
Consider a generalized network $\hat{G}=(V,\hat{E})$ with free arcs $E^{\infty}\subseteq\hat{E}$ and feasible labelings $\hat{\Omega}$. Let $\hat{\Psi}:\hat{\Omega}\to\mathcal{X}$ be the map that takes $l^{\circ}\in\hat{\Omega}$ to the unique equilibrium trajectory $\ell$ satisfying $\ell(0)=l^{\circ}$. Then $\hat{\Psi}$ is a continuous map.
Note that proving the above directly proves Theorem 3.2 by choosing $\hat{G}=G$ and $E^{\infty}=\emptyset$. First, we prove a weaker continuity statement. For every $\theta\in\mathbb{R}_{\geq 0}$, define $\hat{\Psi}_{\theta}\colon\hat{\Omega}\to\mathbb{R}_{\geq 0}^{V}$ by $\hat{\Psi}_{\theta}(l^{\circ})=[\hat{\Psi}(l^{\circ})](\theta)$; that is, we map an initial condition to the value of the resulting equilibrium trajectory at a fixed time $\theta$.
Lemma 5.4.
For any $\theta\in\mathbb{R}_{\geq 0}$, $\hat{\Psi}_{\theta}$ is continuous.
@@
$M\coloneqq\Set{\vartheta\in\mathbb{R}_{\geq 0}}{\hat{\Psi}_{\theta^{\prime}}\text{ is continuous for all }\theta^{\prime}\in[0,\vartheta]}.$
@@
Let $\xi\coloneqq\sup M$ (note that $M\neq\emptyset$ as $0\in M$). As a first step we show that $\xi\in M$. Recall that all equilibrium trajectories are $\kappa$-Lipschitz. This proves that $\hat{\Psi}_{\xi}$ is continuous, because for every $\varepsilon>0$ we can find a $\delta>0$ such that for all $\theta^{\prime}\in[\xi-\frac{\varepsilon}{3\kappa},\xi)$ and $\hat{l}^{\circ}\in B_{\delta}(l^{\circ})\cap\hat{\Omega}$, it holds that
$\|\hat{\Psi}_{\xi}(l^{\circ})-\hat{\Psi}_{\xi}(\hat{l}^{\circ})\|\leq\|\ell(\xi)-\ell(\theta^{\prime})\|+\|\hat{\Psi}_{\theta^{\prime}}(l^{\circ})-\hat{\Psi}_{\theta^{\prime}}(\hat{l}^{\circ})\|+\|\hat{\ell}(\theta^{\prime})-\hat{\ell}(\xi)\|\leq\tfrac{\varepsilon}{3}+\tfrac{\varepsilon}{3}+\tfrac{\varepsilon}{3}\leq\varepsilon.$
@@
Here, $\ell$ and $\hat{\ell}$ are the equilibrium trajectories starting with $l^{\circ}$ and $\hat{l}^{\circ}$, respectively.
@@
We can consider, for every equilibrium trajectory $\ell$, $\ell(\xi)$ as the initial condition of the equilibrium trajectory $\theta^{\prime}\mapsto\ell(\theta^{\prime}+\xi)$. Applying Lemma 5.1, we obtain that for a small duration $[\xi,\xi+\varepsilon]$ the equilibrium trajectory $\ell$ depends continuously on the value of $\ell(\xi)$. But since $\hat{\Psi}_{\xi}$ is continuous, meaning that $\ell(\xi)$ depends continuously on $l^{\circ}$, this shows that $\hat{\Psi}_{\xi+\varepsilon}$ is also continuous, contradicting our choice of $\xi$. ∎
Now we are ready to prove Theorem 5.3.
Proof of Theorem 5.3.
@@
Given an interval $I\subseteq\mathbb{R}_{\geq 0}$, let $\hat{\Psi}_{I}(l^{\circ})$ be the restriction of $\hat{\Psi}(l^{\circ})$ to the time interval $I$, for any $l^{\circ}\in\hat{\Omega}$. It is a basic fact that a sequence of Lipschitz continuous functions that converge pointwise on a compact interval converge uniformly. Thus $\hat{\Psi}_{I}$ is continuous for any compact interval $I$. To extend to the non-compact interval $\mathbb{R}_{\geq 0}$, we again make use of our steady state result. By Theorem 4.3 there is a constant $T$ such that every equilibrium trajectory with starting point in $B_{\delta}(l^{\circ})\cap\hat{\Omega}$ for some fixed $\delta>0$ has reached steady state by time $T$ (note that $T$ depends on distance of $l^{\circ}$ to steady state and $\delta$; more concretely we can choose $T=C_{T}(d(l^{\circ},I_{ss})+\delta)$). Thus for any $l^{\lozenge}\in B_{\delta}(l^{\circ})\cap\hat{\Omega}$ and $\theta>T$, $\hat{\Psi}_{\theta}(l^{\lozenge})=\hat{\Psi}_{T}(l^{\lozenge})+(\theta-T)\cdot\lambda$, where $\lambda$ is the steady-state direction for the generalized network $\hat{G}$, i.e., $\lambda$ is the solution to the thin flow equations for configuration $(\hat{E},E^{\infty})$. Thus $\sup_{\theta\geq 0}\|\hat{\Psi}_{\theta}(l^{\lozenge})-\hat{\Psi}_{\theta}(l^{\circ})\|=\sup_{0\leq\theta\leq T}\|\hat{\Psi}_{\theta}(l^{\lozenge})-\hat{\Psi}_{\theta}(l^{\circ})\|,$ and continuity of $\hat{\Psi}$ follows. ∎
@@ 6 Continuity more generally
First, we prove local continuity, in the sense of Lemma 5.1.
@@
As in the proof of Lemma 5.1, we pick $\varepsilon$ small enough so that we can restrict our attention to the local network at $l^{\circ}$. More precisely, suppose $\varepsilon>0$ is small enough that for all $i$ sufficiently large, the active arcs of $\ell^{(i)}(\theta)$ are in $E^{\prime}_{l^{\circ}}$ for all $\theta\in[0,\varepsilon]$, and all arcs of $E^{*}_{l^{\circ}}$ are resetting arcs of $\ell^{(i)}(\theta)$. As before let $(\lambda,y)$ be a solution to the thin flow equations for the configuration $(E^{\prime}_{l^{\circ}},E^{*}_{l^{\circ}})$ considering transit times ${\tau^{\circ}}$. Note that the steady state direction $\lambda$ is independent of transit times.
Let $\Phi^{(i)}$ be the potential function as defined in (8) for $\ell^{(i)}$ in the network $G^{\tau^{(i)}}$, and let $\Phi^{\circ}$ be the potential for $\ell^{*}$ in $G^{{\tau^{\circ}}}$. Let moreover $\mathsf{OPT}^{(i)}$ be the upper bound on $\Phi^{(i)}$ and $\mathsf{OPT}^{\circ}$ the upper bound on $\Phi^{\circ}$. The dual feasibility set and thus the optimal value depend continuously on $\tau$ and thus $\mathsf{OPT}^{(i)}\to\mathsf{OPT}^{\circ}$ as $i\to\infty$. Since the potential at time 00 depends continuously on transit times and the initial values, we have $\Phi^{(i)}(0)\to\Phi^{\circ}(0)$ as $i\to\infty$. By Lemma 4.12, the time $T_{1}^{(i)}$ after which $\ell^{(i)}$ satisfies the weak steady state properties is upper bounded by $C_{1}(\mathsf{OPT}^{(i)}-\Phi^{(i)}(0))$, where $C_{1}$ depends only on the graph and the capacities and not $\tau$. As $\Phi^{\circ}(0)=\mathsf{OPT}^{\circ}$, we obtain $T_{1}^{(i)}\to 0$ as $i\to\infty$.
Next we show that also the duration after $T_{1}^{(i)}$ until steady state $T_{2}^{(i)}$ goes to 00 as $i\to\infty$. By Lemma 4.14, we have $T_{2}^{(i)}\leq\Delta\max_{v\in V}\gamma^{(i)}_{v}(T_{1}^{(i)})$. The definition of $\Delta$ depends only on the steady-state direction $\lambda$ and is thus independent of the transit times. Since $T_{1}^{(i)}\to 0$, $\ell^{(i)}(T_{1}^{(i)})\to l^{\circ}$; in particular, the distance between $\ell^{(i)}(T_{1}^{(i)})$ and the steady-state set is going to 00. By Lemma 4.13 and Lipschitz continuity of each $\gamma^{(i)}_{v}$ with respect to labels, it follows that $\gamma^{(i)}_{v}(T^{(i)}_{1})\to 0$ for all $v$.
@@
Fix some valid configuration $(E^{\prime},E^{*})$, and let $\lambda^{\nu}$ be the thin flow direction for this configuration in $G^{\nu}$, for any capacity vector $\nu$. Then $\lambda^{\nu}$ depends continuously on $\nu$ in $\mathbb{R}_{>0}^{E}$.
@@
$\sum_{e\in\delta^{+}(v)}x^{\prime}_{e}-\sum_{e\in\delta^{-}(v)}x^{\prime}_{e}$ =$=\begin{cases}u_{0}\;\;\qquad&\text{ if }v=s,\\ 0&\text{ if }v\in V\setminus\{s,t\}.\end{cases}$
This linear system may have multiple solutions. If there exists a solution $(x^{\prime},\ell^{\prime})$ for which $0\leq x^{\prime}_{e}\leq\nu_{e}\ell^{\prime}_{w}$ for all $e=vw$, then it is easy to verify that this satisfies the thin flow conditions; essentially, we have guessed the ordering between label derivatives. Say that $\mathcal{P}$ is a correct (ordered) partition if this holds, and let $N[\mathcal{P}]$ denote the set of capacities for which the linear system for this partition is correct. Note that $N[\mathcal{P}]$ is a closed set. Existence of thin flows implies that for every $\nu$, there is at least one ordered partition $\mathcal{P}$ for which $N[\mathcal{P}]$ contains $\nu$ (there may be more than one; it is possible that for some partition $(V_{1},\ldots,V_{k})$, we obtain a solution with $\ell^{\prime}_{v}=\ell^{\prime}_{w}$ for $v\in V_{j}$, $w\in V_{j+1}$, in which case the partition obtained by merging $V_{j}$ and $V_{j+1}$ would also be correct). Uniqueness of thin flow directions implies that for any correct partition $\mathcal{P}$, and any solution $(x^{\prime},\ell^{\prime})$ to the corresponding system, $\ell^{\prime}=\lambda^{\nu}$.
@@
Let $\lambda^{\nu}$ be the solution to the thin flow equations with configuration $(E^{\prime}_{l^{\circ}},E^{*}_{l^{\circ}})$ in $G^{\nu}$, for any capacities $\nu$. By Lemma 6.2, $\lambda^{\nu}$ depends continuously on $\nu$. For any $\nu$, there exists some $\varepsilon(\nu)>0$ so that $\ell^{\nu}(\theta)=l^{\circ}+\theta X^{\nu}(l^{\circ})$ for $\theta\in[0,\varepsilon(\nu)]$ (this is from considering the equilibrium constructed by the $\alpha$-extension procedure [KS11, CCL15], which we have shown is the only possible equilibrium). Further, if we restrict our attention to some sufficiently small ball around $\nu^{\circ}$, we can choose $\varepsilon(\nu)=\varepsilon$ for some fixed $\varepsilon>0$ for all $\nu$ in this ball. The claim thus follows from the previous lemma. ∎

### v3

@@ 4 Convergence to steady-state
Let $R:=d(l^{\circ},I_{ss})$. We can view the potential $\Phi$ as a function of $l\in\Omega$, by using $l$ rather than $\ell(\theta)$ everywhere in its definition. This function is Lipschitz continuous; denote the Lipschitz constant by $C_{\Phi}$. Then $\mathsf{OPT}-\Phi(0)\leq C_{\Phi}R$. Let $T_{1}=C_{1}(\mathsf{OPT}-\Phi(0))$, with $C_{1}$ as guaranteed by Section 4, so that $\ell$ has the weak steady-state property from time $T_{1}$ onwards. Then $T_{1}\leq C_{1}^{\prime}R$, where $C_{1}^{\prime}=C_{1}C_{\Phi}$.
By Section 4, $\ell$ reaches steady state by time $T_{1}+T_{2}$, where $T_{2}=\Delta\cdot\max_{v\in V}{\gamma_{v}(T_{1})}$. To bound $T_{2}$, we observe that $\gamma_{v}$, viewed as a function of the labeling, is Lipschitz continuous with Lipschitz constant $2|V|$, since the slacks are Lipschitz continuous with Lipschitz constant $2$. Thus $\gamma_{v}(0)\leq 2R|V|$. Since $\|\ell(T_{1})-\ell(0)\|\leq\kappa T_{1}$, this time by Lipschitz continuity of the trajectory, we deduce the very crude bound $\gamma_{v}(T_{1})\leq 2|V|(R+\kappa T_{1})$ for all $v$.
@@ 5 Uniqueness and continuity
[…]
@@
We are now ready to prove uniqueness of equilibrium trajectories.
Proof of Theorem 3.1.
@@
Consider the sequence of equilibrium trajectories $\ell^{(1)},\ell^{(2)},\ell^{(1)},\ell^{(2)},\ldots$ and apply Section 5 to this sequence. Then there exists an $\eta>0$ and a trajectory $\ell^{*}$, so that $\ell^{(1)}(\theta)=\ell^{(2)}(\theta)=\ell^{*}(\theta)$ for $\theta\in[0,\eta]$. This contradicts the assumption that the trajectories $\ell^{(1)}$ and $\ell^{(2)}$ diverge at time 0. ∎
@@
We now prove continuity of equilibrium trajectories.
Proof of Theorem 3.2.
We begin by proving a weaker continuity statement. For every $\theta\in\mathbb{R}_{\geq 0}$, define $\color[rgb]{0,0,0}\Psi_{\theta}\colon\Omega\to\mathbb{R}^{V}$} by ${\color[rgb]{0,0,0}\Psi}_{\theta}(l^{\circ})=[{\color[rgb]{0,0,0}\Psi}(l^{\circ})](\theta)$; that is, we map an initial condition to the value of the resulting equilibrium trajectory at a fixed time $\theta$.
Claim 5.2.
For any $\theta\in\mathbb{R}_{\geq 0}$, ${\color[rgb]{0,0,0}\Psi}_{\theta}$ is continuous.
@@
[…]
@@
Let $\xi\coloneqq\sup M$ (note that $M\neq\emptyset$ as $0\in M$). As a first step we show that $\xi\in M$. Recall that all equilibrium trajectories are $\kappa$-Lipschitz. This proves that ${\color[rgb]{0,0,0}\Psi}_{\xi}$ is continuous, because for every $\varepsilon>0$ we can find a $\delta>0$ such that for $\theta^{\prime}=\xi-\frac{\varepsilon}{3\kappa}$ and ${\color[rgb]{0,0,0}\hat{l}}\in B_{\delta}(l^{\circ})\cap{\color[rgb]{0,0,0}\Omega}$, it holds that
$\|{\color[rgb]{0,0,0}\Psi}_{\xi}(l^{\circ})-{\color[rgb]{0,0,0}\Psi}_{\xi}({\color[rgb]{0,0,0}\hat{l}})\|\leq\|\ell(\xi)-\ell(\theta^{\prime})\|+\|{\color[rgb]{0,0,0}\Psi}_{\theta^{\prime}}(l^{\circ})-{\color[rgb]{0,0,0}\Psi}_{\theta^{\prime}}({\color[rgb]{0,0,0}\hat{l}})\|+\|\hat{\ell}(\theta^{\prime})-\hat{\ell}(\xi)\|\leq\tfrac{\varepsilon}{3}+\tfrac{\varepsilon}{3}+\tfrac{\varepsilon}{3}\leq\varepsilon.$
@@
Here, $\ell$ and $\hat{\ell}$ are the equilibrium trajectories starting with $l^{\circ}$ and $^{\color[rgb]{0,0,0}\hat{l}}$, respectively.
@@
We can consider, for every equilibrium trajectory $\ell$, $\ell(\xi)$ as the initial condition of the equilibrium trajectory $\theta^{\prime}\mapsto\ell(\theta^{\prime}+\xi)$. Applying Section 5, we obtain that for a small duration $\xi,\xi+\varepsilon]$ the equilibrium trajectory $\ell$ depends continuously on the value of $\ell(\xi)$. But since ${\color[rgb]{0,0,0}\Psi}_{\xi}$ is continuous, meaning that $\ell(\xi)$ depends continuously on $l^{\circ}$, this shows that ${\color[rgb]{0,0,0}\Psi}_{\xi+\varepsilon}$ is also continuous, contradicting our choice of $\xi$. ∎
@@
Given an interval $I\subseteq\mathbb{R}_{\geq 0}$, let $\color[rgb]{0,0,0}\Psi_{I}(l)$} be the restriction of $\color[rgb]{0,0,0}\Psi(l)$} to the time interval $I$, for any ${\color[rgb]{0,0,0}l}\in\Omega$. It is a basic fact that a sequence of Lipschitz continuous functions with a common Lipschitz constant that converge pointwise on a compact interval converge uniformly. Thus ${\color[rgb]{0,0,0}\Psi}_{I}$ is continuous for any compact interval $I$. To extend to the non-compact interval $\mathbb{R}_{\geq 0}$, we again make use of our steady state result. By Theorem 4.3 there is a constant $T$ such that every equilibrium trajectory with starting point in $B_{\delta}(l^{\circ})\cap\Omega$ for some fixed $\delta>0$ has reached steady state by time $T$ (note that $T$ depends on distance of $l^{\circ}$ to steady state and $\delta$; more concretely we can choose $T=C_{T}(d(l^{\circ},I_{ss})+\delta)$). Thus for any ${\color[rgb]{0,0,0}l}\in B_{\delta}(l^{\circ})\cap\Omega$ and $\theta>T$, ${\color[rgb]{0,0,0}\Psi}_{\theta}({\color[rgb]{0,0,0}l})={\color[rgb]{0,0,0}\Psi}_{T}({\color[rgb]{0,0,0}l})+(\theta-T)\cdot\lambda$, where $\lambda$ is the steady-state direction, i.e., $\lambda$ is the thin flow direction for configuration $(E,E^{\infty})$. Thus $\sup_{\theta\geq 0}\|{\color[rgb]{0,0,0}\Psi}_{\theta}({\color[rgb]{0,0,0}l})-{\color[rgb]{0,0,0}\Psi}_{\theta}(l^{\circ})\|=\sup_{0\leq\theta\leq T}\|{\color[rgb]{0,0,0}\Psi}_{\theta}({\color[rgb]{0,0,0}l})-{\color[rgb]{0,0,0}\Psi}_{\theta}(l^{\circ})\|,$ and continuity of $\color[rgb]{0,0,0}\Psi$} follows. ∎
Piecewise-constant inflow rates.
The results on continuity over a compact interval and uniqueness translate easily to the setting of piecewise-constant inflow rates. Simply apply the continuity and uniqueness results to each maximal interval for which the inflow is constant; since the composition of continuous functions is continuous, there are no difficulties. More formally, local continuity as stated in Section 5 still holds, where the inflow rate is taken to be the constant inflow over some small interval of time starting from a time at interest, and so Theorem 3.1 as well as Section 5 (continuity over a compact interval) still follow. For a constant inflow, we were able to show continuity over the unbounded interval $\mathbb{R}_{\geq 0}$; this relied on the convergence to steady-state, which certainly does not hold in general for a time-varying inflow. So this result does not extend, unless the inflow is constant after some finite time.
@@ 6 Continuity more generally
First, we prove local continuity, in the sense of Section 5.
@@
As in the proof of Section 5, we pick $\varepsilon$ small enough so that we can restrict our attention to the local network at $l^{\circ}$. More precisely, suppose $\varepsilon>0$ is small enough that for all $i$ sufficiently large, the active arcs of $\ell^{(i)}(\theta)$ are in $E^{\prime}_{l^{\circ}}$ for all $\theta\in[0,\varepsilon]$, and all arcs of $E^{*}_{l^{\circ}}$ are resetting arcs of $\ell^{(i)}(\theta)$. As before let $(\lambda,y)$ be a solution to the thin flow equations for the configuration $(E^{\prime}_{l^{\circ}},E^{*}_{l^{\circ}})$ considering transit times $\tau^{\circ}$}. Note that the steady state direction $\lambda$ is independent of transit times.
Let $\Phi^{(i)}$ be the potential function as defined in (8) for $\ell^{(i)}$ in the network $G^{\tau^{(i)}}$, and let $\Phi^{\circ}$ be the potential for $\ell^{*}$ in $G^{{\tau^{\circ}}}$. Let moreover $\mathsf{OPT}^{(i)}$ be the upper bound on $\Phi^{(i)}$ and $\mathsf{OPT}^{\circ}$ the upper bound on $\Phi^{\circ}$. The dual feasibility set and thus the optimal value depend continuously on $\tau$ and thus $\mathsf{OPT}^{(i)}\to\mathsf{OPT}^{\circ}$ as $i\to\infty$. Since the potential at time 0 depends continuously on transit times and the initial values, we have $\Phi^{(i)}(0)\to\Phi^{\circ}(0)$ as $i\to\infty$. By Section 4, the time $T_{1}^{(i)}$ after which $\ell^{(i)}$ satisfies the weak steady state properties is upper bounded by $C_{1}(\mathsf{OPT}^{(i)}-\Phi^{(i)}(0))$, where $C_{1}$ depends only on the graph and the capacities and not $\tau$. As $\Phi^{\circ}(0)=\mathsf{OPT}^{\circ}$, we obtain $T_{1}^{(i)}\to 0$ as $i\to\infty$.
Next we show that also the duration after $T_{1}^{(i)}$ until steady state $T_{2}^{(i)}$ goes to 0 as $i\to\infty$. By Section 4, we have $T_{2}^{(i)}\leq\Delta\max_{v\in V}\gamma^{(i)}_{v}(T_{1}^{(i)})$. The definition of $\Delta$ depends only on the steady-state direction $\lambda$ and is thus independent of the transit times. Since $T_{1}^{(i)}\to 0$, $\ell^{(i)}(T_{1}^{(i)})\to l^{\circ}$; in particular, the distance between $\ell^{(i)}(T_{1}^{(i)})$ and the steady-state set is going to 0. By Section 4 and Lipschitz continuity of each $\gamma^{(i)}_{v}$ with respect to labels, it follows that $\gamma^{(i)}_{v}(T^{(i)}_{1})\to 0$ for all $v$.
@@
Fix some valid configuration $(E^{\prime},E^{*})$, and let $\lambda^{\nu}$ be the thin flow direction for this configuration in $G^{\nu}$, for any capacity vector $\nu$. Then $\lambda^{\nu}$ depends continuously on $\nu$.
@@
$\sum_{e\in\delta^{+}(v)}x^{\prime}_{e}-\sum_{e\in\delta^{-}(v)}x^{\prime}_{e}$ $=\begin{cases}u_{0}\;\;\qquad&\text{ if }v=s,\\ 0&\text{ if }v\in V\setminus\{s,t\}.\end{cases}$
This linear system may have multiple solutions. If there exists a solution $(\ell^{\prime},x^{\prime})$ for which $0\leq x^{\prime}_{e}\leq\nu_{e}\ell^{\prime}_{w}$ for all $e=vw$, then it is easy to verify that this satisfies the thin flow conditions; essentially, we have guessed the ordering between label derivatives. Say that $\mathcal{P}$ is a correct (ordered) partition if this holds, and let $N[\mathcal{P}]$ denote the set of capacities for which the linear system for this partition is correct. Note that $N[\mathcal{P}]$ is a closed set. Existence of thin flows implies that for every $\nu$, there is at least one ordered partition $\mathcal{P}$ for which $N[\mathcal{P}]$ contains $\nu$ (there may be more than one; it is possible that for some partition $(V_{1},\ldots,V_{k})$, we obtain a solution with $\ell^{\prime}_{v}=\ell^{\prime}_{w}$ for $v\in V_{j}$, $w\in V_{j+1}$, in which case the partition obtained by merging $V_{j}$ and $V_{j+1}$ would also be correct). Uniqueness of thin flows implies that for any correct partition $\mathcal{P}$, and any solution $(\ell^{\prime},x^{\prime})$ to the corresponding system, $\ell^{\prime}=\lambda^{\nu}$.
@@
Let $\lambda^{\nu}$ be the solution to the thin flow equations with configuration $(E^{\prime}_{l^{\circ}},E^{*}_{l^{\circ}})$ in $G^{\nu}$, for any capacities $\nu$. By Section 6, $\lambda^{\nu}$ depends continuously on $\nu$. For any $\nu$, there exists some $\varepsilon(\nu)>0$ so that $\ell^{\nu}(\theta)=l^{\circ}+\theta X^{\nu}(l^{\circ})$ for $\theta\in[0,\varepsilon(\nu)]$ (this is from considering the equilibrium constructed by the $\alpha$-extension procedure [KS11, CCL15], which we have shown is the only possible equilibrium). Further, if we restrict our attention to some sufficiently small ball around $\nu^{\circ}$, we can choose $\varepsilon(\nu)=\varepsilon$ for some fixed $\varepsilon>0$ for all $\nu$ in this ball. The claim thus follows from the previous lemma. ∎